OpenStudy (anonymous):

e^(x) - e^(-x) - 10 = 0 Having some trouble solving for x on this one.

OpenStudy (anonymous):

can be verified why here: http://en.wikipedia.org/wiki/Hyperbolic_function \[\frac{ e^x - e^{-x} }{ 2 } = \sinh(x)\]

OpenStudy (anonymous):

your calculator should have the sinh^{-1} function

OpenStudy (experimentx):

make use of inverse hyperbolic function

OpenStudy (anonymous):

This question is from a chapter on logarithmic and exponential functions--before the hyperbolic function is introduced. Is that the only way to solve for x?

OpenStudy (anonymous):

Alternatively, you can treat this as a quadratic: \[e^x-\frac{1}{e^x}-10=0\\ e^{2x}-1-10e^x=0\] Substitute \(y=e^x\): \[y^2-10y-1=0\]

OpenStudy (anonymous):

Yeah, that is more what I was looking for

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

So you're multiplying both sides by e^x? Ok--I see it now. Thanks.